# A Survey on the Safety and Security Threats of Computer-Using Agents: JARVIS or Ultron?

### v2

@@ 5.4.2 LLM-as-a-judge Measurements
LLM-based measurement leverages LLMs, such as GPT-4, to compute evaluation metrics based on natural language understanding, reasoning, and contextual judgment. Unlike rule-based methods that rely on fixed logic, LLM-based approaches utilize the interpretive abilities of LLMs to handle complex and open-ended scenarios, making them ideal for tasks where deterministic rules fall short.
This approach is increasingly adopted in recent agent safety and capability benchmarks. For example, R-Judge (Yuan et al., 2024) uses an LLM-as-a-judge framework to score open-ended safety analyses, while TrustAgent (Hua et al., 2024) employs GPT-4 to assess both helpfulness and safety in agent outputs. PrivacyLens (Shao et al., 2024) applies a GPT-based few-shot classifier to determine whether sensitive information is inferable from an agent’s action.
Similarly, BrowserART (Kumar et al., 2024) and AgentHarm (Andriushchenko et al., 2024) use GPT-4o to classify harmful behaviors and evaluate refusals. CASA (Qiu et al., 2025) adopts GPT-4o across metrics to assess cultural and social awareness, while ASB (Zhang et al., 2024b) uses LLMs to evaluate whether agents properly refuse unsafe instructions.
LLM-based methods are highly flexible and capable of capturing nuanced behaviors and contextual subtleties that rule-based systems often miss. However, they may suffer from variability across model versions, increased computational cost, and potential inconsistencies in subjective judgments.
@@ 5.4.3 Manual Judge Measurements
Nevertheless, manual labels remain a valuable source of ground truth. For instance, R-Judge (Yuan et al., 2024) incorporates a human-labeled test set to assess the quality of LLM-generated safety analyses, using manual annotations as the gold standard to validate automated or LLM-based scoring methods.
@@ 6.1 Key Insights
  • UI Understanding Difficulties: Current benchmarking efforts reveal that many CUAs demonstrate suboptimal safety performance, reflecting gaps in robustness and risk awareness (Zhang et al., 2024e; Andriushchenko et al., 2024; Lee et al., 2024a; Zhang et al., 2024b). A significant portion of this shortfall arises from immature grounding techniques, which hinder agents’ ability to reliably interpret multimodal perceptions, especially vision tasks (Zheng et al., 2024; Zhang and Zhang, 2023; Liu et al., 2025). These shortcomings underscore the need for more holistic training and test scenarios that address diverse threat models.
  • Limited Experimental Scenarios: Many CUAs are tested in highly constrained settings that fail to capture the breadth of real-world tasks. For instance, the action space is restricted by excluding multi-touch or irregular gestures in (Zhang et al., 2023), while (Zhang and Zhang, 2023) focuses on small-scale GUI agents in a single environment (AITW), limiting broader investigations. Additionally, PC-Agent has only explored productivity scenarios, leaving potential social and entertainment applications largely unexamined (Liu et al., 2025).
@@ 6.2 Future Directions
  • Integrated Defense Mechanisms: Research on robust defenses spans active adversarial mitigation, environmental verification, and backdoor detection. Proposed methods include integrating modules for trustworthiness checks and leveraging multi-agent approaches for role-specific security tasks (Chen et al., 2025b; Zeng et al., 2024; Tian et al., 2023).
@@ Appendix A Complete Taxonomy
for tree= forked edges, draw, rounded corners, node options=align=center, s sep=12pt, calign=center, grow=east, reversed=true, parent anchor=east, child anchor=north, font=, , where level=1text width=100pt, fill=blue!10, rotate=90, where level=2text width=100pt, fill=pink!30, child anchor=west, where level=3text width=50pt, fill=yellow!30, child anchor=west, where level=4text width=100pt, fill=cyan!20, child anchor=west, where level=5child anchor=west, [Computer-Using Agent Safety, fill=gray!20, rotate=90 [Threats [Intrinsic Threats [Perception [UI Understanding and Grounding Difficulties, text width=150pt [Chen et al. (2025a), Pahuja et al. (2025), Nong et al. (2024), Zheng et al. (2024), Liu et al. (2025) , text width = 250pt] ] ] [Brain [Scheduling Errors, text width=150pt [Zhang and Zhang (2023), Deng et al. (2024) ] ] [Misalignment, text width=150pt [Ma et al. (2024) ] ] [Hallucination, text width=150pt [Deng et al. (2024) ] ] [Excessive Context Length, text width=150pt [Zhang and Zhang (2023), Yang et al. (2024a) ] ] [Social and Cultural Concerns, text width=150pt [Qiu et al. (2025) ] ] [Response Latency, text width=150pt [Wen et al. (2023), Zhang and Zhang (2023), Li et al. (2020), Li et al. (2021) ] ] ] [Action [API Call Error, text width=150pt [Deng et al. (2024), Nong et al. (2024), Zhang and Zhang (2023) ] ] ] ] [Extrinsic Threats [Adversarial Attack, text width=150pt, fill=cyan!20 [Wu et al. (2024a), Ma et al. (2024), Zhang et al. (2024c) , text width=300pt, fill=white] ] [Prompt Injection Attack, text width=150pt, fill=cyan!20 [Zhan et al. (2024), Liao et al. (2024), Wu et al. (2024b), Xu et al. (2024), Debenedetti et al. (2024) , text width=400pt, fill=white] ] [Jailbreak, text width =150pt, fill=cyan!20 [Mo et al. (2024), Chu et al. (2024), Zou et al. (2023), Liu et al. (2023a), Kumar et al. (2024), Tian et al. (2023), Zhang et al. (2024d) , text width=400pt, fill=white] ] [Backdoor Attack, text width =150pt, fill=cyan!20 [Yang et al. (2024b), Wang et al. (2024), Zhu et al. (2025), Chen et al. (2024b) , text width = 400pt, fill=white] ] [Reasoning Gap Attack, text width = 150pt, fill=cyan!20 [Chen et al. (2025b) , fill=white] ] [System Sabotage, text width =150pt, fill=cyan!20 [Luo et al. (2025) , fill=white] ] [Web Hacking, text width =150pt, fill=cyan!20 [(Fang et al., 2024b), fill=white] ] ] ] [Defenses [Environmental Constraints, text width=210pt, fill=cyan!20 [Yang et al. (2024c), Nong et al. (2024), Zhang et al. (2023), text width=300pt, fill=white] ] [Input Validation, text width=210pt, fill=cyan!20 [Kumar et al. (2024), Wen et al. (2023), Zhang et al. (2024c), text width=300pt, fill=white] ] [Defensive Prompting, text width=210pt, fill=cyan!20 [Debenedetti et al. (2024), Zhang et al. (2024c), Wu et al. (2024a), Chen et al. (2025b), text width=450pt, fill=white] ] [Data Sanitization,text width=210pt, fill=cyan!20 [Yang et al. (2024b),text width=100pt, fill=white] ] [Adversarial Training,text width=210pt, fill=cyan!20 [Wu et al. (2024a),text width=100pt, fill=white] ] [Output Monitoring, text width=210pt, fill=cyan!20 [Fang et al. (2024a),Wen et al. (2023), Hua et al. (2024), text width=300pt, fill=white] ] [Model Inspection, text width=210pt, fill=cyan!20 [Wang et al. (2025), Yang et al. (2024b), Huang et al. (2025), Zhu et al. (2025), text width=400pt, fill=white] ] [Cross Verification, text width=210pt, fill=cyan!20 [Zeng et al. (2024), Yang et al. (2024a), Luo et al. (2025), text width=300pt, fill=white] ] [Continuous Learning and Adaptation, text width=210pt, fill=cyan!20 [Tian et al. (2023), Ma et al. (2024), Luo et al. (2025), text width=300pt, fill=white] ] [Transparentize, text width=210pt, fill=cyan!20 [Hu et al. (2024), Sager et al. (2025), text width=200pt, fill=white] ] [Topology-Guided, text width=210pt, fill=cyan!20 [Wang et al. (2025), text width=100pt, fill=white] ] [Perception Algorithms Synergy, text width=210pt, fill=cyan!20 [Zheng et al. (2024), Nong et al. (2024), Liu et al. (2025), Yang et al. (2024a), text width=400pt, fill=white] ] [Planning-Centric Architecture Refinement, text width=210pt, fill=cyan!20 [Zhang and Zhang (2023), Deng et al. (2024), Wen et al. (2023), Liu et al. (2025), text width=420pt, fill=white] ] [Pre-defined Regulatory Compliance, text width=210pt, fill=cyan!20 [Zhang et al. (2024e), Mahadevan et al. (2025), Chen et al. (2025c), text width=350pt, fill=white] ] ] ]

### v3

@@ 5.4.2 LLM-as-a-judge Measurements
LLM-based measurement leverages LLMs, such as general models like GPT-4 or fine-tuned models, to compute evaluation metrics based on natural language understanding, reasoning, and contextual judgment (Luo et al., 2025a). Unlike rule-based methods that rely on fixed logic, LLM-based approaches utilize the interpretive abilities of LLMs to handle complex and open-ended scenarios, making them ideal for tasks where deterministic rules fall short.
This approach is increasingly adopted in recent agent safety and capability benchmarks. For example, R-Judge (Yuan et al., 2024) uses an LLM-as-a-judge framework to score open-ended safety analyses, while TrustAgent (Hua et al., 2024) employs GPT-4 to assess both helpfulness and safety in agent outputs. ToolEmu (Ruan et al., 2023) uses automatic LLM evaluators to score safety and helpfulness for each trajectory. Both PrivacyLens (Shao et al., 2024) and AgentDAM (Zharmagambetov et al., 2025) applies a LLM-based evaluator to determine whether sensitive information is inferable from an agent’s action or output.
Similarly, BrowserART (Kumar et al., 2024) and AgentHarm (Andriushchenko et al., 2024) use GPT-4o to classify harmful behaviors and evaluate refusals. CASA (Qiu et al., 2025) adopts GPT-4o across metrics to assess cultural and social awareness, SafeArena (Tur et al., 2025) feeds GPT-4o each agent’s trajectory and metadata to assign one of the four ARIA risk levels, ASB (Zhang et al., 2024b) uses LLMs to evaluate whether agents properly refuse unsafe instructions, and MLA-Trust (Yang et al., 2025c) employs auto-classifiers to evaluate response toxicity and the misguided rate. Furthermore, OpenAgentSafety (Vijayvargiya et al., 2025) uses GPT-4.1 to label each trajectory into one of four predefined safety categories to capture unsafe intent that may not manifest in the final environment state. OS-Harm (Kuntz et al., 2025) also employs an LLM judge to decide task completion, label safety, and pinpoint the first unsafe step, with human annotations validating and confirming the LLM judge’s effectiveness. HAICOSYSTEM scores trajectories with LLM judges using scenario-specific checklists, assigning five risk dimensions on a 10-10 to 0 scale, plus Goal Completion and Tool-Use Efficiency on a 0 to $10$ scale.
Furthermore, RedTeamCUA (Liao et al., 2025), RiOSWorld (Yang et al., 2025b) and WASP (Evtimov et al., 2025) rely on LLM to identify trajectories where the agent shows intent to pursue the unsafe objective even without completing it. Rather than using a single LLM judge, VPI-Bench (Cao et al., 2025) employs a majority vote of three frontier models to decide whether a malicious instruction was attempted and whether it was completed. DoomArena framework (Boisvert et al., 2025) also supports plugging in LLM-based monitors to review intermediate agent thoughts and actions, flagging subtler breaches that its rule-based scripts might miss.
LLM-based methods are highly flexible and capable of capturing nuanced behaviors and contextual subtleties that rule-based systems often miss. However, they may suffer from variability across model versions, increased computational cost, and potential inconsistencies or bias in subjective judgments.
@@ 5.4.3 Manual Judge Measurements
Nevertheless, manual labels remain a valuable source of ground truth. For instance, R-Judge (Yuan et al., 2024) incorporates a human-labeled test set to assess the quality of LLM-generated safety analyses, using manual annotations as the gold standard to validate automated or LLM-based scoring methods. Similarly, ToolEmu (Ruan et al., 2023) employ human annotators to label emulation quality and agent safety/helpfulness, providing a reference set to validate the LLM judges. In addition, SafeArena (Tur et al., 2025) conduct manual assessments of each trajectory using the ARIA framework.
@@ 6.1 Key Insights
  • UI Understanding Difficulties: Current benchmarking efforts reveal that many CUAs demonstrate suboptimal safety performance, reflecting gaps in robustness and risk awareness (Zhang et al., 2024e; Andriushchenko et al., 2024; Lee et al., 2024a; Zhang et al., 2024b). A significant portion of this shortfall arises from immature grounding techniques, which hinder agents’ ability to reliably interpret multimodal perceptions, especially vision tasks (Zheng et al., 2024; Zhang and Zhang, 2023; Liu et al., 2025b). These shortcomings underscore the need for more holistic training and test scenarios that address diverse threat models.
  • Limited Experimental Scenarios: Many CUAs are tested in highly constrained settings that fail to capture the breadth of real-world tasks. For instance, the action space is restricted by excluding multi-touch or irregular gestures in (Zhang et al., 2023), while (Zhang and Zhang, 2023) focuses on small-scale GUI agents in a single environment (AITW), limiting broader investigations. Additionally, PC-Agent has only explored productivity scenarios, leaving potential social and entertainment applications largely unexamined (Liu et al., 2025b).
@@ 6.2 Future Directions
  • Integrated Defense Mechanisms: Research on robust defenses spans active adversarial mitigation, environmental verification, and backdoor detection. Proposed methods include integrating modules for trustworthiness checks and leveraging multi-agent approaches for role-specific security tasks (Chen et al., 2025d; Zeng et al., 2024; Tian et al., 2023).
@@ Appendix A Complete Taxonomy
for tree= forked edges, draw, rounded corners, node options=align=center, s sep=12pt, calign=center, grow=east, reversed=true, parent anchor=east, child anchor=north, font=, , where level=1text width=100pt, fill=blue!10, rotate=90, where level=2text width=100pt, fill=pink!30, child anchor=west, where level=3text width=50pt, fill=yellow!30, child anchor=west, where level=4text width=100pt, fill=cyan!20, child anchor=west, where level=5child anchor=west, [Computer-Using Agent Safety, fill=gray!20, rotate=90 [Threats [Intrinsic Threats [Perception [UI Understanding and Grounding Difficulties, text width=150pt [Chen et al. (2025c), Pahuja et al. (2025), Nong et al. (2024), Zheng et al. (2024), Liu et al. (2025b) , text width = 350pt] ] ] [Brain [Scheduling Errors, text width=150pt [Zhang and Zhang (2023), Deng et al. (2024) ] ] [Misalignment, text width=150pt [Ma et al. (2024) ] ] [Hallucination, text width=150pt [Deng et al. (2024) ] ] [Excessive Context Length, text width=150pt [Zhang and Zhang (2023), Yang et al. (2024a) ] ] [Social and Cultural Concerns, text width=150pt [Qiu et al. (2025) ] ] [Response Latency, text width=150pt [Wen et al. (2023), Zhang and Zhang (2023), Li et al. (2020), Li et al. (2021) ] ] ] [Action [API Call Error, text width=150pt [Deng et al. (2024), Nong et al. (2024), Zhang and Zhang (2023) ] ] ] ] [Extrinsic Threats [Adversarial Attack, text width=150pt, fill=cyan!20 [Wu et al. (2024a), Ma et al. (2024), Zhang et al. (2024c), Aichberger et al. (2025), Zhao et al. (2025), Wu et al. (2025) , text width=300pt, fill=white] ] [Prompt Injection Attack, text width=150pt, fill=cyan!20 [Zhan et al. (2024), Liao et al. (2024), Wu et al. (2024b), Xu et al. (2024), Debenedetti et al. (2024), Mudryi et al. (2025), Liu et al. (2023b), Lupinacci et al. (2025), Kuntz et al. (2025), Liao et al. (2025), Evtimov et al. (2025), Wu et al. (2025), Cao et al. (2025), Shapira et al. (2025), Chen et al. (2025a), Liu et al. (2025a), Wang et al. (2025a), Nakash et al. (2024), Johnson et al. (2025), Wang et al. (2025e), Zhan et al. (2025), Lu et al. (2025), Wang et al. (2025g), Wang et al. (2025c) , text width=400pt, fill=white] ] [Jailbreak, text width =150pt, fill=cyan!20 [Mo et al. (2024), Chu et al. (2024), Zou et al. (2023), Liu et al. (2023a), Kumar et al. (2024), Tian et al. (2023), Zhang et al. (2024d), Mao et al. (2025), Kuntz et al. (2025), Gu et al. (2024), Qi et al. (2025) , text width=450pt, fill=white] ] [Memory injection Attack, text width =150pt, fill=cyan!20 [Patlan et al. (2025a), Patlan et al. (2025b), text width=400pt, fill=white] ] [Backdoor Attack, text width =150pt, fill=cyan!20 [Yang et al. (2024b), Wang et al. (2024), Zhu et al. (2025b), Chen et al. (2024b), Boisvert et al. , Ye et al. (2025), Wang et al. (2025f), Cheng et al. (2025), Lupinacci et al. (2025) , text width = 400pt, fill=white] ] [Reasoning Gap Attack, text width = 150pt, fill=cyan!20 [Chen et al. (2025d) , fill=white] ] [System Sabotage, text width =150pt, fill=cyan!20 [Luo et al. (2025b) , fill=white] ] [Web Hacking, text width =150pt, fill=cyan!20 [(Fang et al., 2024b), fill=white] ] ] ] [Defenses [Environmental Constraints, text width=210pt, fill=cyan!20 [Yang et al. (2024c), Nong et al. (2024), Zhang et al. (2023), Mahadevan et al. (2025), Huang et al. (2025), text width=400pt, fill=white] ] [Input Validation, text width=210pt, fill=cyan!20 [Kumar et al. (2024), Wen et al. (2023), Zhang et al. (2024c), Tshimula et al. (2024), Shi et al. (2025a), Zhong et al. (2025), Ferrag et al. (2025), text width=450pt, fill=white] ] [Defensive Prompting, text width=210pt, fill=cyan!20 [Debenedetti et al. (2024), Zhang et al. (2024c), Wu et al. (2024a), Chen et al. (2025d), text width=450pt, fill=white] ] [Data Sanitization,text width=210pt, fill=cyan!20 [Yang et al. (2024b), Jones et al. (2025), Wang et al. (2025b),text width=300pt, fill=white] ] [Adversarial Training,text width=210pt, fill=cyan!20 [ Yu et al. (2020), Wu et al. (2024a), Zhou et al. (2025), Yu et al. (2025),text width=400pt, fill=white] ] [Output Monitoring, text width=210pt, fill=cyan!20 [Fang et al. (2024a),Wen et al. (2023), Hua et al. (2024), Lee et al. (2025), Shi et al. (2025b), text width=500pt, fill=white] ] [Model Inspection, text width=210pt, fill=cyan!20 [Wang et al. (2025d), Yang et al. (2024b), Huang et al. (2025), Zhu et al. (2025b), text width=400pt, fill=white] ] [Cross Verification, text width=210pt, fill=cyan!20 [Zeng et al. (2024), Huang et al. (2024), Yang et al. (2024a), Xiang et al. (2024), Luo et al. (2025b), Barua et al. (2025), Zhu et al. (2025a), Li et al. (2025), Fan and Li (2025), text width=500pt, fill=white] ] [Continuous Learning and Adaptation, text width=210pt, fill=cyan!20 [Tian et al. (2023), Ma et al. (2024), Luo et al. (2025b), Zhan et al. (2025), Zhang et al. (2025b), text width=500pt, fill=white] ] [Transparentize, text width=210pt, fill=cyan!20 [Hu et al. (2024), Sager et al. (2025), Chen et al. (2025b), text width=300pt, fill=white] ] [Topology-Guided, text width=210pt, fill=cyan!20 [Wang et al. (2025d), text width=100pt, fill=white] ] [Perception Algorithms Synergy, text width=210pt, fill=cyan!20 [Zheng et al. (2024), Nong et al. (2024), Liu et al. (2025b), Yang et al. (2024a), text width=400pt, fill=white] ] [Planning-Centric Architecture Refinement, text width=210pt, fill=cyan!20 [Zhang and Zhang (2023), Deng et al. (2024), Wen et al. (2023), Liu et al. (2025b), text width=420pt, fill=white] ] [Pre-defined Regulatory Compliance, text width=210pt, fill=cyan!20 [Zhang et al. (2024e), Mahadevan et al. (2025), Chen et al. (2025e), Zhang et al. (2025a), text width=450pt, fill=white] ] ] ]